### v1

# A Symmetry-Preserving Reduced-Order Observer

Abstract
A symmetry-preserving, reduced-order state observer is presented for the unmeasured part of a system’s state, where the nonlinear system dynamics exhibit symmetry under the action of a Lie group. The proposed observer takes advantage of this symmetry through the use of a moving frame that constructs invariant mappings of the measurements. Sufficient conditions for the observer to be asymptotically stable are developed by studying the stability of an invariant error system. As an illustrative example, the observer is applied to the problem of rigid-body velocity estimation, which demonstrate how exploiting the symmetry of the system can simplify the stabilization of the estimation error dynamics.
I INTRODUCTION
Methods for designing state observers for nonlinear systems are limited and there are no general techniques that guarantee global convergence of the estimation error, as there are in the linear case [1, Ch. 15]. Provably effective state estimation strategies are inevitably limited to special classes of systems. Here, we leverage symmetries in a dynamical system’s structure to aid observer design and stability analysis. These so-called symmetry-preserving observers presented by Bonnabel et al. in [2] and [3] are adapted to systems whose dynamics are invariant under the action of a Lie group. The idea is to design an observer that is also invariant, i.e., for which the observer dynamics also preserve this Lie group symmetry. This approach allows the observer’s convergence properties to be analyzed more easily because of simplifications afforded by the system symmetry.
Existing approaches to symmetry-preserving observers only consider the full-order case, however, in which the entire state of the system is estimated ([2, 3, 4]). In many scenarios, part of the system’s state may be known with negligible error or may be obtained as the output of an observer whose design is independent of the rest of the system’s state. For example, attitude observers for aircraft or spacecraft often do not rely on the rigid body’s translational dynamics ([5, 6]). Another example is the problem of wind estimation from aircraft motion ([7, 8, 9]), where the main goal is to obtain estimates of wind and air-relative velocity – not to re-estimate the aircraft’s position, attitude, and angular velocity. More generally, the problem of disturbance estimation falls into this category where the internal state of the system is known but the disturbance is not. In these scenarios, reduced-order observers, in the sense of Karagiannis et al. in [10] and [11], are of particular interest where only the unmeasured part of the system’s state is estimated. The aim of observer design is to render a particular set, characterized by zero state estimation error, positively invariant and globally asymptotically attractive. Here, we focus on reduced-order observers that are also symmetry-preserving.
The remainder of this paper is organized as follows. Section II introduces the preliminary concepts that will be used in the development of a reduced-order, symmetry-preserving pre-observer in Section III. Next, sufficient conditions for the pre-observer to be an asymptotically stable observer are presented in Section IV. Finally, the main results are applied to the example of rigid-body velocity estimation in Section V followed by concluding remarks in Section VI.
II PRELIMINARIES
II-A Transformation Groups, the Moving Frame, and Invariant Dynamics
II-A1 Transformation Groups [12, 13]
Consider a differentiable (i.e., $C^{\infty}$ or smooth) manifold $\mathcal{X}$ on which a Lie group $G$ acts via the mapping
${\varphi}:G\times\mathcal{X}\to\mathcal{X},~{}(g,{x})\mapsto{\varphi}_{g}({x})$
such that (i) the identity element $e$ in $G$ induces the identity transformation ${\varphi}_{e}({x})={x}$ for all ${x}\in\mathcal{X}$, and (ii)  the composition of group actions satisfies ${\varphi}_{g}\circ{\varphi}_{h}={\varphi}_{g\ast h}$, where $\circ$” denotes the composition of mappings and $\ast$” is group multiplication. The inverse transformation ${\varphi}_{g}^{-1}$ is given by the action of the inverse group element – i.e., ${\varphi}_{g}^{-1}={\varphi}_{g^{-1}}$. The Lie group $G$ is said to act freely on $\mathcal{X}$ if ${\varphi}_{g}({x})={x}$ implies $g$ is the identity element, $e$. The collection $\{{\varphi}_{g}\}_{g\in G}$ is called a transformation group. The $g$-orbit of a point ${x}\in\mathcal{X}$ is the set $\{{\varphi}_{g}({x})~{}|~{}g\in G\}$.
II-A2 The Moving Frame [2, 13, 14]
A moving frame is a mapping $\gamma:\mathcal{X}\to G$ that has the following equivariance property (illustated in Fig. 1):
$\gamma({\varphi}_{g}({x}))\ast g=\gamma({x})$ (1)
[…]
Figure 1: Equivariance of the moving frame $\gamma$ and its construction via the cross-section $\mathcal{K}$
It may be associated with a coordinate cross-section $\mathcal{K}$ that transversely intersects $g$-orbits on $\mathcal{X}$. Informally, for an $r$-dimensional Lie group $G$ acting freely on the $n$-dimensional manifold $\mathcal{X}$, let $\varphi_{g}^{\mathrm{inv}}$ be the part of $\varphi_{g}$ that maps to an $r$-dimensional submanifold of $\mathcal{X}$ such that it is invertible with respect to $g$ in a neighborhood of the identity element $e\in G$. Then, one can select a constant ${k}$ in the image of $\varphi_{g}^{\mathrm{inv}}$ that defines the unique point at which the $g$-orbit of a generic point ${x}$ intersects the $(n-r)$-dimensional cross-section $\mathcal{K}$. In other words, it is obtained by solving the normalization equation
$\varphi_{h}^{\mathrm{inv}}({x})={k}$
for $h\in G$. The local solution $h=\gamma({x})$ defines the moving frame.
II-A3 Invariant Dynamics [2, 12]
Consider the dynamical control system
$\dot{{x}}={f}({x},{u})$ (2)
where ${x}(t)\in\mathcal{X}$ (a differentiable manifold), ${u}(t)\in\mathcal{U}$ (a set), and ${f}(\cdot,{u}):\mathcal{X}\to\mathsf{T}_{x}\mathcal{X}$ (a vector field on $\mathcal{X}$ for each ${u}\in\mathcal{U}$). Here, $\mathsf{T}_{x}\mathcal{X}$ denotes the tangent space to $\mathcal{X}$ at ${x}$. Let $\{({\varphi}_{g}({x}),{\psi}_{g}({u}))\}_{g\in G}$ be a transformation group on $\mathcal{X}\times\mathcal{U}$. The mapping ${\varphi}_{g}:\mathcal{X}\to\mathcal{X}$ induces the tangent map $\mathrm{T}{\varphi}_{g}({x}):\mathsf{T}_{x}\mathcal{X}\to\mathsf{T}_{{\varphi}_{g}({x})}\mathcal{X}$ at ${x}$. Note if $\mathcal{X}=\mathbb{R}^{n}$, then $\mathrm{T}{\varphi}_{g}({x})$ is simply the Jacobian matrix, $\partial{\varphi}_{g}({x})/\partial{x}$. The system (2) is called $G$-invariant if
${f}({\varphi}_{g}({x}),{\psi}_{g}({u}))=\mathrm{T}{\varphi}_{g}({x})\cdot{f}({x},{u})$
Here, $\cdot$” denotes the application of the tangent map to a tangent vector. It follows that the tangent map of ${\varphi}_{g\ast h}({x})={\varphi}_{g}({\varphi}_{h}({x}))$ satisfies
$\mathrm{T}{\varphi}_{g\ast h}({x})=\mathrm{T}{\varphi}_{g}({\varphi}_{h}({x}))\cdot\mathrm{T}{\varphi}_{h}({x})$
A function $I:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is called an invariant if $I({\varphi}_{g}({x}),{\psi}_{g}({u}))=I(x,u)$ for all $g\in G$. Suppose $G$ acts freely on $\mathcal{X}$. Then, there locally exist $n-r$ functionally independent invariants $(I_{1}({x}),\dots,I_{n-r}({x}))$.
II-B Immersion and Invariance Observers [10, 11]
Consider a dynamical system whose state is described by a measured part, ${y}\in\mathcal{Y}\subseteq\mathbb{R}^{p}$, and an unmeasured part, ${x}\in\mathcal{X}\subseteq\mathbb{R}^{n}$, with dynamics
$\dot{{x}}$ $={f}({x},{y})$
$\dot{{y}}$ $={h}({x},{y})$
The dynamical system
$\dot{{z}}={\alpha}({z},{y})$
where ${z}\in\mathbb{R}^{q(\geq n)}$, is called a (global) reduced-order observer for ${x}$ if there exists a smooth manifold
$\mathcal{Z}=\left\{({x},{y},{z})~{}|~{}{\theta}({z},{y})={\phi}({x},{y})\right\}$ (3)
defined by mappings ${\theta}$ and ${\phi}$ that are left invertible with respect to their first argument such that $\mathcal{Z}$ is positively invariant and (globally) asymptotically attractive. The estimate of ${x}$ is then given by
$\hat{{x}}={\phi}^{(\mathrm{L},\cdot)}({\theta}({z},{y}),{y})$
where ${\phi}^{(\mathrm{L},\cdot)}$ denotes the functional left inverse of ${\phi}$ with respect to its first argument (i.e., ${\phi}^{(\mathrm{L},\cdot)}({\phi}({x},{y}),{y})={x}$).
III INVARIANT PRE-OBSERVER
Consider a system whose state is given by a measured part, ${y}\in\mathcal{Y}$, and an unmeasured part, ${x}\in\mathcal{X}\subseteq\mathbb{R}^{n}$. Here, $\mathcal{Y}$ is a $p$-dimensional differentiable manifold and $\mathcal{X}$ is an open subset of $\mathbb{R}^{n}$ that contains the origin. The dynamics of this system are given by
$\dot{{x}}$ $={f}({x},{y},{u})$ (4a)
$\dot{{y}}$ $={h}({x},{y},{u})$ (4b)
where ${u}\in\mathcal{U}$ is the known “input” to the system. It is not necessarily just composed of control inputs, but rather is a known signal on which a particular Lie group acts. Here, the dynamics of the measured part of the state, ${y}$, may be expressed intrinsically, that is, without specifying a local coordinate chart.
We consider systems of the form (4) that are invariant under actions of some Lie group, $G$.
Assumption 1
The system (4) is $G$-invariant under the transformation group $\{({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))\}_{g\in G}$, where $G$ is an $r$-dimensional Lie group. That is,
$\mathrm{T}{\varphi}_{g}({x})\cdot{f}({x},{y},{u})$ $={f}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
$\mathrm{T}{\varrho}_{g}({y})\cdot{h}({x},{y},{u})$ $={h}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
Furthermore, ${\varphi}_{g}({x})$ is linear in ${x}$.
We can now describe what it means for a reduced-order pre-observer in the sense of [10, 11] to be symmetry-preserving under the transformation group considered in Assumption 1. Briefly, a pre-observer is an observer for which there is not (yet) any claim about error convergence. We postulate the form of an observer for the unmeasured part of the state, ${x}$, that preserves invariance of the state estimate dynamics. Inspired by [2], consider the following definition.
Definition 1 ($G$-invariant reduced-order pre-observer)
The dynamical system
$\dot{{z}}={\alpha}({z},{y},{u})$ (5)
with output
$\hat{{x}}={z}+{\beta}({y})$ (6)
for some smooth map ${\beta}:\mathcal{Y}\to\mathcal{X}$ is a $G$-invariant reduced-order pre-observer if the system
$\dot{\hat{{x}}}={\alpha}(\hat{{x}}-{\beta}({y}),{y},{u})+\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$ (7)
is $G$-invariant and the manifold
$\mathcal{Z}=\left\{({z},{x},{y})\in\mathcal{X}\times\mathcal{X}\times\mathcal{Y}~{}|~{}{z}={x}-{\beta}({y})\right\}$ (8)
is positively invariant. A $G$-invariant pre-observer is a $G$-invariant observer if $\mathcal{Z}$ is asymptotically attractive.
This prescription of the zero-error manifold $\mathcal{Z}$ is not quite as general as the case described in [11]. We instead consider the condition that defines $\mathcal{Z}$ to be linear in ${z}$ and ${x}$ (and for ${z}$ to be the same dimension as ${x}$). In other words, we choose ${\theta}({z},{y})={z}$ and ${\phi}({x},{y})={x}-{\beta}({y})$ in Eq. (3). This simplification reveals an intuitive choice for ${\beta}$ in the following lemma based on the underlying geometry.
Lemma 1
Suppose there exists a moving frame $\gamma:\mathcal{Y}\to G$ that only depends on ${y}\in\mathcal{Y}$, and let ${\ell}:\mathcal{Y}\to\mathcal{X}$ be a smooth map. If
${\beta}({y})={\varphi}_{\gamma({y})^{-1}}\Big{(}{\ell}\big{(}{\varrho}_{\gamma({y})}({y})\big{)}\Big{)}$ (9)
then the following commutative identities (illustrated in Fig. 2) hold for all $g\in G$ and ${y}\in\mathcal{Y}$:
${\beta}({\varrho}_{g}({y}))={\varphi}_{g}({\beta}({y}))$ (10)
$\mathrm{T}{\beta}({\varrho}_{g}({y}))\cdot\mathrm{T}{\varrho}_{g}({y})=\mathrm{T}{\varphi}_{g}({\beta}({y}))\cdot\mathrm{T}{\beta}({y})$ (11)
In other words, ${\beta}$ commutes with the transformation group.
[…]
Figure 2: Commutative relationship between ${\beta}$ and the transformation group
Proof:
Beginning with the definition of ${\beta}$, we have
${\beta}({\varrho}_{g}({y}))={\varphi}_{\gamma({\varrho}_{g}({y}))^{-1}}\Big{(}{\ell}\big{(}{\varrho}_{\gamma({\varrho}_{g}({y}))}({\varrho}_{g}({y}))\big{)}\Big{)}$
By the equivariance property (1) of the moving frame,
${\beta}({\varrho}_{g}({y}))={\varphi}_{(\gamma\ast g^{-1})^{-1}}\Big{(}{\ell}\big{(}{\varrho}_{\gamma\ast g^{-1}}({\varrho}_{g}({y}))\big{)}\Big{)}$
Using the composition properties for group elements and group actions,
${\beta}({\varrho}_{g}({y}))$ $={\varphi}_{g\ast\gamma^{-1}}\Big{(}{\ell}\big{(}{\varrho}_{\gamma\ast g^{-1}\ast g}({y})\big{)}\Big{)}$
$={\varphi}_{g}\Big{(}{\varphi}_{\gamma^{-1}}\big{(}{\ell}\big{(}{\varrho}_{\gamma}({y})\big{)}\big{)}\Big{)}$
$={\varphi}_{g}({\beta}({y}))$
Finally, (11) follows directly from the properties of the tangent map. ∎
Using Lemma 1, a $G$-invariant pre-observer is readily constructed for a system satisfying Assumption 1.
Theorem 1
Suppose Assumption 1 and the conditions of Lemma 1 hold. Let the vector field ${\alpha}(\cdot,{y},{u}):\mathcal{X}\to\mathsf{T}\mathcal{X}$ be defined by
${\alpha}({z},{y},{u})={f}({z}+{\beta}({y}),{y},{u})-\mathrm{T}{\beta}({y})\cdot{h}({z}+{\beta}({y}),{y},{u})$ (12)
Then, the dynamical system
$\dot{{z}}={\alpha}({z},{y},{u})$ (13)
with output
$\hat{{x}}={z}+{\beta}({y})$ (14)
is a $G$-invariant, reduced-order pre-observer.
Proof:
We begin by showing invariance of the state estimate dynamics (7). Define
${F}(\hat{{x}},{x},{y},{u})={\alpha}(\hat{{x}}-{\beta}({y}),{y},{u})+\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$
Then,
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})=\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{f}(\hat{{x}},{y},{u})\\ -\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot\mathrm{T}{\beta}({y})\cdot\left({h}(\hat{{x}},{y},{u})-{h}({x},{y},{u})\right)$
The assumed linearity of ${\varphi}_{g}$ implies $\mathrm{T}{\varphi}_{g}({x})$ does not depend on the choice of base point ${x}$. Therefore, $\mathrm{T}{\varphi}_{g}(\hat{{x}})=\mathrm{T}{\varphi}_{g}({\beta}({y}))$, and Lemma 1 can be used along with the invariance of ${f}$ to obtain
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})={f}({\varphi}_{g}(\hat{{x}}),{\varrho}_{g}({y}),{\psi}_{g}({u}))\\ -\mathrm{T}{\beta}\left({\varrho}_{g}({y})\right))\cdot\mathrm{T}{\varrho}_{g}({y})\cdot\left({h}(\hat{{x}},{y},{u})-{h}({x},{y},{u})\right)$
Since ${h}$ is also $G$-invariant, we have
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})={f}({\varphi}_{g}(\hat{{x}}),{\varrho}_{g}({y}),{\psi}_{g}({u}))\\ -\mathrm{T}{\beta}\left({\varrho}_{g}({y})\right)\cdot\Big{(}{h}({\varphi}_{g}(\hat{{x}}),{\varrho}_{g}({y}),{\psi}_{g}({u}))\\ -{h}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))\Big{)}$
By Lemma 1, we recognize
${\varphi}_{g}(\hat{{x}})={\varphi}_{g}({z})+{\beta}({\varrho}_{g}({y}))$
Then, it follows that
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})={\alpha}({\varphi}_{g}(\hat{{x}})-{\beta}({\varrho}_{g}({y})),{\varrho}_{g}({y}),{\psi}_{g}({u}))\\ +\mathrm{T}{\beta}\left({\varrho}_{g}({y})\right)\cdot{h}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
Therefore,
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})={F}({\varphi}_{g}(\hat{{x}}),{\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
That is, the system (12)-(14) is $G$-invariant. Next, we show the zero error manifold $\mathcal{Z}$ given in (8) is positively invariant. Since ${z}-{x}+{\beta}({y})={0}$ on $\mathcal{Z}$, we verify that
${\alpha}({x}-{\beta}({y}),{y},{u})-{f}({x},{y},{u})+\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$
$\quad={f}({x},{y},{u})-\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$
$\qquad-{f}({x},{y},{u})+\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$
$\quad={0}$
Thus, referring to (13), trajectories originating in $\mathcal{Z}$ remain in $\mathcal{Z}$. It follows that (12)-(14) is a $G$-invariant reduced-order pre-observer. ∎
IV INVARIANT OBSERVER
We now aim to find sufficient conditions for the pre-observer in Theorem 1 to be a $G$-invariant reduced-order observer. That is, we seek conditions under which $\mathcal{Z}$ is asymptotically attractive. Like [2], we consider error coordinates that are $G$-invariant. Specifically, we take
${\eta}({z},{x},{y})={\varphi}_{\gamma({y})}({z})+{\ell}({\varrho}_{\gamma({y})}({y}))-{\varphi}_{\gamma({y})}({x})$ (15)
to be invariant coordinates that are non-zero if and only if $({z},{x},{y})\notin\mathcal{Z}$. Thus, ${\eta}\to{0}$ as $t\to\infty$ implies $\mathcal{Z}$ is asymptotically attractive. Let ${X}={\varphi}_{\gamma({y})}({x})$, ${Y}={\varrho}_{\gamma({y})}({y})$, and ${U}={\psi}_{\gamma({y})}({u})$. Using the moving frame to define these transformed points means $({X},{Y},{U})$ comprises a complete set of invariants [13, Ch. 8]. As will be shown shortly, the stability of the pre-observer (13) depends only ${\eta}$ and the invariants ${X}$, ${Y}$, and ${Z}$ (see Remark 1).
To derive sufficient conditions for asymptotic stability, we will make use of the following result.
Lemma 2
Let ${\lambda}:\mathcal{Y}\to\mathcal{X}$ be the map
${\lambda}({y};{x})={\varphi}_{\gamma({y})}({x})$
where ${x}$ is held constant. Then,
$\mathrm{T}{\lambda}({y};{x})\cdot{h}({x},{y},{u})=\mathrm{T}{\lambda}({Y};{X})\cdot{h}({X},{Y},{U})$ (16)
The following proof of Lemma 2 is illustrated in Fig. 3.
[…]
Figure 3: Invariance of ${\lambda}$ and its tangent map
Proof:
First, we recognize ${\lambda}$ is invariant since
${\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))$ $={\varphi}_{\gamma({\varrho}_{g}({y}))}({\varphi}_{g}({x}))$
$={\varphi}_{\gamma({y})\ast g^{-1}}({\varphi}_{g}({x}))$
$={\varphi}_{\gamma({y})}({x})$
$={\lambda}({y};{x})$
where we again use the equivariance of the moving frame $\gamma$. Let
${\Lambda}({y};{x})={\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))$
Since ${\Lambda}$ is a composition of maps, it follows that its tangent map is
$\mathrm{T}{\Lambda}({y};{x})=\mathrm{T}{\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))\cdot\mathrm{T}{\varrho}_{g}({y})$
Then since ${\Lambda}({y};{x})={\lambda}({y};{x})$, we have
$\mathrm{T}{\lambda}({y};{x})=\mathrm{T}{\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))\cdot\mathrm{T}{\varrho}_{g}({y})$
Applying $\mathrm{T}{\lambda}({y};{x})$ to the vector field ${h}$, we obtain
$\mathrm{T}{\lambda}({y};{x})\cdot{h}({x},{y},{u})\\ =\mathrm{T}{\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))\cdot\mathrm{T}{\varrho}_{g}({y})\cdot{h}({x},{y},{u})$
Finally, by the $G$-invariance of the dynamics,
$\mathrm{T}{\lambda}({y};{x})\cdot{h}({x},{y},{u})\\ =\mathrm{T}{\lambda}({\varrho}_{g}({y});{\varphi}_{g}({x}))\cdot{h}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
which verifies Eq. (16). ∎
Finally, sufficient conditions for (13) to be a $G$-invariant reduced-order observer are given as follows.
Theorem 2
Suppose the assumptions of Theorem 1 hold. The $G$-invariant pre-observer (13) is a $G$-invariant observer if the origin ${\eta}={0}$ of the invariant error system
$\dot{{\eta}}={f}({X}+{\eta},{Y},{U})-{f}({X},{Y},{U})\\ -\mathrm{T}{\beta}({Y})\cdot\big{(}{h}({X}+{\eta},{Y},{U})-{h}({X},{Y},{U})\big{)}\\ +\mathrm{T}{\lambda}({Y};{\eta})\cdot{h}({X},{Y},{U})$ (17)
is asymptotically stable uniformly in ${X}$, ${Y}$, and ${U}$.
Proof:
By definition, the pre-observer (13) is an observer if the zero error manifold $\mathcal{Z}$ is positively invariant and asymptotically attractive or, equivalently, if the state estimation error dynamics have a globally asymptotically stable equilibrium at the origin. It remains for us to show that the estimation error dynamics are given by the invariant error system (17). Since ${\varphi}_{g}({x})$ is linear in ${x}$, we can write
${\eta}={\varphi}_{\gamma({y})}({z}+{\beta}({y}))-{\varphi}_{\gamma({y})}({x})$
Thus, the time derivative of ${\eta}$ satisfies
$\dot{{\eta}}=\mathrm{T}{\varphi}_{\gamma}({z}+{\beta}({y}))\cdot\big{(}{\alpha}({z},{y},{u})+\mathrm{T}{\beta}({y}))\cdot{h}({x},{y},{u})\big{)}\\ -\mathrm{T}{\varphi}_{\gamma}({x})\cdot{f}({x},{y},{u})+\mathrm{T}{\lambda}({y};{z}+{\beta}({y})-x)\cdot{h}({x},{y},{u})$
Substituting the definition of ${\alpha}$ from Theorem 1 and again using the linearity of ${\varphi}_{g}(\cdot)$, we have
$\dot{{\eta}}=\mathrm{T}{\varphi}_{\gamma}({z}+{\beta}({y}))\cdot\big{(}{f}({z}+{\beta}({y}),{y},{u})\big{)}-\mathrm{T}{\varphi}_{\gamma}({x})\cdot{f}({x},{y},{u})\\ -\mathrm{T}{\varphi}_{\gamma}({\beta}({y}))\cdot\mathrm{T}{\beta}({y})\cdot\big{(}{h}({z}+{\beta}({y}),{y},{u})-{h}({x},{y},{u})\big{)}\\ +\mathrm{T}{\lambda}({y};{z}+{\beta}({y})-x)\cdot{h}({x},{y},{u})$
Applying the invariance of ${f}$ and ${h}$ through the use of Lemma 1 yields
$\dot{{\eta}}={f}\big{(}{\varphi}_{\gamma}({z}+{\beta}({y})),{Y},{U}\big{)}-{f}\big{(}{X},{Y},{U}\big{)}\\ -\mathrm{T}{\beta}({Y})\cdot\Big{(}{h}\big{(}{\varphi}_{\gamma}({z}+{\beta}({y})),{Y},{U}\big{)}-{h}\big{(}{X},{Y},{U}\big{)}\Big{)}\\ +\mathrm{T}{\lambda}({y};{z}+{\beta}({y})-x)\cdot{h}({x},{y},{u})$
Substituting
${z}={\varphi}_{\gamma^{-1}}({\eta})-{\beta}({y})+{x}$
and using Lemma 2, we obtain Eq. (17). ∎
Remark 1
The error system (17) depends only on the invariant error ${\eta}$ and the invariants ${X}$, ${Y}$, and ${U}$, which can be reduced to a set of $n+p-r$ functionally independent invariants, ${I}({x},{y},{u})$ [13, Ch. 8]. This observation is consistent with the full-order case considered in [2, Theorem 3].
V EXAMPLE: RIGID-BODY VELOCITY OBSERVER
Consider a rigid aircraft instrumented with an accelerometer, gyroscope, magnetometer, and GNSS receiver such that its position, ${q}$, and attitude rotation matrix, ${R}_{\mathrm{IB}}$, are known without error. Furthermore, assume the angular velocity, ${\omega}$, and body-frame specific force, ${a}$, (obtained from filtered accelerometer readings) are available as inputs for the observer design. However, suppose that the body velocity, ${v}=(\mathrm{u,v,w})$ is not directly measured. The aim is to design a reduced-order velocity observer for the system
$\begin{gathered}\underbrace{\dot{{v}}}_{\dot{{x}}}=\underbrace{{v}\times{\omega}+{R}_{\mathrm{IB}}^{\mathsf{T}}{g}+{a}}_{{f}({x},{y},{u})}\\ \underbrace{\begin{pmatrix}\dot{{q}}\\ \dot{{R}}_{\mathrm{IB}}\end{pmatrix}}_{\dot{{y}}}=\underbrace{\begin{pmatrix}{R}_{\mathrm{IB}}{v}\\ {R}_{\mathrm{IB}}{S}({\omega})\end{pmatrix}}_{{h}({x},{y},{u})}\end{gathered}$ (18)
where ${S}(\cdot)$ is the skew-symmetric cross product equivalent matrix satisfying ${S}({a}){b}={a}\times{b}$ for 3-vectors ${a}$ and ${b}$ and ${g}$ is the gravity vector.
Proposition 1
The system (18) is $\mathsf{SO}(3)$-invariant with respect to the transformation group
${\varphi}_{g}({x})={R}_{g}{v},\quad{\varrho}_{g}({y})=\begin{pmatrix}{q}\\ {R}_{\mathrm{IB}}{R}_{g}^{\mathsf{T}}\end{pmatrix},\quad{\psi}_{g}({u})=\begin{pmatrix}{R}_{g}{\omega}\\ {R}_{g}{a}\end{pmatrix}$
where ${R}_{g}\in G=\mathsf{SO}(3)$.
Proof:
We have
$\mathrm{T}{\varphi}_{g}({x})\cdot{f}({x},{y},{u})$ $={R}_{g}({v}\times{\omega})+{R}_{g}{R}_{\mathrm{IB}}^{\mathsf{T}}{g}+{R}_{g}{a}$
$={R}_{g}{v}\times{R}_{g}{\omega}+({R}_{\mathrm{IB}}{R}_{g}^{\mathsf{T}})^{\mathsf{T}}{g}+{R}_{g}{a}$
$={f}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
and
$\mathrm{T}{\varphi}_{g}({x})\cdot{h}({x},{y},{u})$ $=\begin{pmatrix}{R}_{\mathrm{IB}}{v}\\ {R}_{\mathrm{IB}}{S}({\omega}){R}_{g}^{\mathsf{T}}\end{pmatrix}$
$=\begin{pmatrix}{R}_{\mathrm{IB}}{R}_{g}^{\mathsf{T}}{R}_{g}{v}\\ {R}_{\mathrm{IB}}{R}_{g}^{\mathsf{T}}{S}({R}_{g}{\omega})\end{pmatrix}$
$={h}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
Here, we have used the property that ${S}({R}{\xi})={R}{S}({\xi}){R}^{\mathsf{T}}$ for any ${R}\in\mathsf{SO}(3)$ and ${\xi}\in\mathbb{R}^{3}$. ∎
Since ${R}_{\mathrm{IB}}$ is an element of the Lie group $G$, the moving frame is simply
$\gamma({y})={R}_{\mathrm{IB}}$
Because the transformation group is also linear in the measured part of the state, we can choose ${\ell}$ to simply be
${\ell}({y})={L}{q}$
where ${L}\in\mathbb{R}^{3\times 3}$. Therefore,
${\beta}({y})={R}_{\mathrm{IB}}^{\mathsf{T}}{L}{q}$
Applying Theorem 1, we have
${\alpha}({z},{x},{y})=\underbrace{({z}+{R}_{\mathrm{IB}}^{\mathsf{T}}{L}{q})\times{\omega}+{R}_{\mathrm{IB}}^{\mathsf{T}}{g}+{a}}_{{f}({z}+{\beta}({y}),{y},{u})}\\ \underbrace{+{S}({\omega}){R}_{\mathrm{IB}}^{\mathsf{T}}{L}{q}-{R}_{\mathrm{IB}}^{\mathsf{T}}{L}{R}_{\mathrm{IB}}({z}+{R}_{\mathrm{IB}}^{\mathsf{T}}{L}{q})}_{-\mathrm{T}{\beta}({y})\cdot{h}({z}+{\beta}({y}),{y},{u})}$
with the estimate of ${v}$ given by
$\hat{{v}}={z}+{R}_{\mathrm{IB}}^{\mathsf{T}}{L}{q}$
The sufficient condition given in Theorem 2 reduces to the requirement that the system
$\dot{{\eta}}=-{L}{\eta}$
is asymptotically stable with respect to $\mathcal{Z}$. Therefore, if $(-{L})$ is Hurwitz, then the pre-observer $\dot{{z}}={\alpha}({z},{y},{u})$ is a globally exponentially stable, reduced-order, $\mathsf{SO}(3)$-invariant observer.
As a numerical example, consider the maneuvering flight trajectory shown in Fig. 4. For the initial condition $\hat{{v}}(0)={0}$ and gain matrix ${L}=10\mathbb{I}$, the time history of velocity estimates is shown in Fig. 5. To stress the observer, we include noisy measurements of ${y}$ and ${u}$. Specifically, suppose
$y_{q}$ $={q}+{w}_{q}$ $u_{\omega}$ $={\omega}+{w}_{\omega}$
$y_{R_{\mathrm{IB}}}$ $={R}_{\mathrm{IB}}\exp({S}({w}_{R_{\mathrm{IB}}}))$ $u_{a}$ $={a}+{w}_{a}$
where ${w}_{q}$, ${w}_{R_{\mathrm{IB}}}$, ${w}_{\omega}$, and ${w}_{a}$ are zero-mean, Gaussian, continuous-time, “white noise” with power spectral densities $5\times 10^{-4}\mathbb{I}~{}\frac{\mathrm{m}^{2}}{\mathrm{Hz}}$, $10^{-7}\mathbb{I}~{}\frac{1}{\mathrm{Hz}}$, $10^{-5}\mathbb{I}~{}\frac{(\mathrm{rad}/\mathrm{s})^{2}}{\mathrm{Hz}}$, and $2\times 10^{-2}\mathbb{I}~{}\frac{(\mathrm{m}/\mathrm{s}^{2})^{2}}{\mathrm{Hz}}$, respectively. Fig. 6 shows the velocity estimates when ${y}$ and ${u}$ are corrupted by a realization of these random processes. While proof of stability for this case is beyond the scope of this paper, the results shown in Fig. 6 are indicative of the observer’s inherent robustness to disturbances, as expected from the fact that the undisturbed invariant error system is globally exponentially stable.
Figure 4: Maneuvering aircraft
Figure 5: Velocity estimates
Figure 6: Velocity estimates with noisy measurements and inputs
VI CONCLUSIONS
A new symmetry-preserving, reduced-order observer has been presented. This approach is beneficial when part of the system’s state is known with negligible error, avoiding unnecessary re-estimation of known signals and reducing computational complexity. Furthermore, the observer preserves symmetry. That is, the state estimate dynamics are invariant under the action of some Lie Group. As a result, the state estimate error dynamics are also invariant under the group action. Exploiting these symmetries can simplify the selection of observer parameters as seen with the example of a rigid-body velocity observer. Tuning the proposed observer consists of finding a smooth map ${\beta}$ such that the origin of the invariant error system is asymptotically stable. For some systems, this problem is reduced to choosing a gain matrix, ${L}$, as shown in the example. By leveraging the geometry of the problem, the proposed observer simplifies both the design process and stability analysis, providing a powerful tool for nonlinear systems.